If (x) represents the fractional part of x then 5^200÷8 is
 

Grade:11

1 Answers

Aditya Gupta
5 years ago
5^200= 25^100= 8k+1, k is an integer
So 5^200/8= k+1/8
Hence fractional part of the above term is 1/8.